# Saliency-Guided DETR for Moment Retrieval and Highlight Detection

### v1

@@ 4.3.1 Motivation
The QVHighlights dataset contains only 10k training samples. Previous studies [16, 19] have shown that this dataset size is insufficient for training a model to effectively address a complex task like Moment Retrieval. As a result, even pretraining on a noisy dataset can significantly improve the model’s metrics. For instance, pretraining framework developed in [19] boosted the model performance from 35.47 to 43.63 mAP@avg. Nevertheless, the metrics in the zero-shot mode remained relatively low at only 10.87 mAP@avg [19], leading us to believe that creating a dataset for pretraining remains an unsolved challenge.
@@ 4.3.2 InterVid-MR Dataset
Recent advancements in video-language modeling have been driven by the annotation of large scale, high-quality datasets like InternVid [41]. Unlike VideoCC [26] and other datasets previously used for pretraining [11, 16], InternVid was annotated using LLMs and does not rely on pseudo-labeling, resulting in significantly less annotation noise. To construct our dataset, we used the subset InternVid-10M-FLT [41] derived from the original InternVid by sampling clips with UMT-SIM scores [18] ranking among the top 30 percent to ensure high quality.
Unfortunately, InternVid-10M-FLT cannot be directly used to train MR models because each caption is linked to only one video segment in the annotations, whereas in practice multiple video segments may closely match a given query. We have made several steps to address this issue and developed a new dataset named InterVid-MR, which is suitable for Moment Retrieval and Highlight Detection tasks.
First, we additionally filtered the captions of the InternVid-10M-FLT dataset to remove meaningless ones. Next, to align the data with the QVHighlights dataset, we selected video segments no longer than 150 seconds for each caption from the InternVid-10M-FLT dataset. These segments were created by trimming the original videos around the time intervals associated with the captions.
@@
$S=\text{L2\_Norm}(F_{v})\cdot\text{L2\_Norm}(F_{c})^{T},$ (16)
where $F_{v}$ represents the embedding of the video clip, and $F_{c}$ represents the caption embedding.
@@
After annotating the Moment Retrieval task, we proceeded to annotate the Highlight Detection task. We used the similarity scores calculated in the previous step to assign each clip within the identified positive intervals a grade ranging from 1 to 4. The grades are calculated using the following formula:
@@
where $S$ is the score from the positive interval to be graded, $\mu$ is the mean score of the positive clips, $\sigma$ is the standard deviation of the positive clips’ scores, and $S_{\text{max}}$ is the maximum score of the positive clips.
As a result, we obtain a dataset that can be used to solve both Moment Retrieval and Highlight Detection tasks simultaneously.
@@ 4.3.3 Impact of the Pretraining
Using our data annotation framework, we collected a dataset consisting of 150k samples. To ensure the effectiveness of the proposed data annotation strategy for solving Moment Retrieval and Highlight Detection tasks, we measured the zero-shot performance of a model trained on different amounts of pre-train data on the QVHighlights dataset. As shown in Figure  2, even 20k samples in the training dataset are sufficient to achieve an MR mAP@avg above 45. Using all the data available, we surpassed the 50 MR mAP@avg threshold, which significantly exceeds the results of the pretraining strategy proposed in [19].
30303030606060609090909012012012012015015015015045454545505050505555555560606060Number of Pretrain PointsmAP@avg (%)QVHighlightsPretrainFinetune
@@ 4.4 Training Details
We train SG-DETR using two attention-pooling layers in the Local Saliency Head and two cross-attention transformer layers in the SGCA Encoder. The model architecture includes three layers of encoders and three layers of decoders. The hidden dimension of the transformer layers is set to 256. we employ the AdamW optimizer with a weight decay of 1e-1. A dropout rate of 0.1 is applied throughout the model, with a higher dropout rate of 0.5 in the projection layers. Our query selector mechanism generates 25 queries per instance. Unless otherwise specified, the following loss balancing parameters are used: $\lambda_{\text{L1}}=10$, $\lambda_{\text{gIoU}}=1$, $\lambda_{\text{CE}}=5$, $\lambda_{\text{Centerness}}=1$, $\lambda_{\text{detr}}=1$, $\lambda_{\text{atss}}=1$, $\lambda_{\text{hl}}=1$, $\lambda_{\text{aux}}=1$.
@@
  • QVHighlight: The model is trained for 150 epochs with a batch size of 128 and a learning rate of 5e-4. We employ the WarmupMultiStepLR scheduler, utilizing 45 warmup steps, with two milestones set at 100 and 125 epochs. The learning rate is decayed by a factor of 0.5 at each milestone.
  • Charades-STA: The model is trained for 100 epochs with a batch size of 64 and a learning rate of 5e-4. Video features are extracted from 1-second clips. During training, binary highlight scores were generated based on whether a clip overlapped more than halfway with a relevant moment retrieval interval
  • TACoS: The model is trained for 120 epochs with a batch size of 32 and a learning rate of 5e-4. Video features are extracted from 2-second clips. Binary highlight scores were generated the same way as in Charades-STA.
Table 3: Experimental results on the QVHighlights test set. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training. † denotes using audio modality. Bold letters indicate the best results.
@@
Ours 72.2 56.6 73.2 55.8 54.1 43.76 69.13
Ours ZS 63.9 49.6 67.5 49.0 48.3 43.0 68.0
Ours w/ PT 74.2 60.4 76.2 60.8 58.8 44.7 71.0
Table 4: Experimental results on the Charades-STA and TACoS test sets. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training. † denotes using audio modality. Bold letters indicate the best results.
@@
Ours 70.2 49.5 59.1 44.7 29.9 40.9
Ours w/ PT 71.1 52.8 60.7 46.4 33.9 42.4
@@ 4.5 Comparison with the state-of-the-art
To evaluate the effectiveness of SG-DETR in jointly addressing Moment Retrieval and Highlight Detection tasks, we conducted experiments on the QVHighlights test split, as shown in Table 3. Remarkably, even without pretraining, SG-DETR significantly outperforms all existing methods that do not leverage LLMs, and it delivers performance comparable to the MrBLIP model despite the latter having considerably more parameters. After pretraining on our constructed InterVid-MR dataset, SG-DETR achieves competitive zero-shot performance, with 48.3 mAP and 68.0 HIT@1 results. When fine-tuned on the QVHighlights dataset, the pre-trained SG-DETR surpasses all existing methods, including those based on LLMs, with 58.8 mAP and 71.0 HIT@1 results.
Moment Retrieval
We use the Charades-STA and TACoS datasets as benchmarks to evaluate the performance of SG-DETR exclusively in the moment retrieval task. The results are presented in Table 4. SG-DETR achieves state-of-the-art performance on these benchmarks, both with and without leveraging additional data from our pretraining framework. The results highlight the model’s robustness and ability to generalize effectively across varying task conditions.
4.6 Ablation study
To evaluate the contribution of each component in our approach, we conducted a comprehensive ablation study, the results of which are shown in Table 2. The baseline model (a) consists of a DETR-based architecture [16, 19], augmented with the local saliency head described in Sec. 3.2 and employing standard cross-attention as the text-to-vision interaction mechanism [25]. In experiment (b), we assess the effect of incorporating the SGCA encoder, as detailed in Sec. 3.3. It increases the speed of convergence of the model and pushes model performance forward, especially the moment retrieval part. Experiment (c) introduces the IoU scoring mechanism, described in Sec. 3.6, resulting in further growth of moment retrieval performance. Next, in experiment (d), we investigate the influence of the GSM module, discussed in Sec. 3.4 on model performance. It sufficiently improves not only moment retrieval performance but especially the quality of HD. Finally, in experiment (e), we replace the standard DETR head with a hybrid ATSS-DETR head, as outlined in Sec. 3.5, drastically improving MR performance.
5 Conclusions
We introduced the SG-DETR, a unified framework for Moment Retrieval and Highlight Detection, and presented the new InterVid-MR dataset for pretraining. Evaluations on benchmarks such as QVHighlights, Charades-STA and TACoS demonstrate that the SG-DETR sets new state-of-the-art results across all tasks. Our ablation studies highlight the effectiveness of key components like the Saliency-Guided Cross-Attention encoder and hybrid ATSS-DETR head, offering a scalable and robust solution for video-language modeling.
6 Limitations
Moment Retrieval and Highlight Detection tasks present significant challenges due to the ambiguity of textual queries and the subjective nature of video content. While SG-DETR achieves state-of-the-art performance, it may face difficulties with abstract or nuanced queries where the boundaries of relevant video segments are less clearly defined.
7 Acknowledgements
We want to express our sincere gratitude to Mikhail Chernyshov for his valuable insights and constructive discussions, which significantly contributed to shaping the direction of the work.
Supplementary Material
A Results on Highlight Detection Benchmarks
To evaluate the effectiveness of SG-DETR in addressing the highlight detection task independently, we evaluate its performance on the Youtube-hl [38] and TVSum [35] benchmarks. YouTube Highlights consists of 433 videos divided into six categories, and TVSum consists of 50 videos divided into ten categories. The results are presented in Table 5 and Table 6. The SG-DETR model achieves state-of-the-art performance on the YouTube-HL benchmark when additional data is utilized. Furthermore, SG-DETR achieves highly competitive results even in the zero-shot setting. The TVSum dataset contains only a limited number of videos for training in each domain, which appears to be insufficient for our model. Although the model achieves state-of-the-art performance specifically in the VT, PK, and BT domains, it overall performs slightly worse than TR-DETR. Additionally, due to the instability of training on such a small dataset, fine-tuning the model with additional data does not yield any improvements.
Table 5: Experimental results of mAP on the Youtube-hl. † denotes using audio modality. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training, Bold letters indicate the best results.
@@ A Results on Highlight Detection Benchmarks
Ours 75.1 79.7 76.9 71.0 74.5 83.2 76.7
Ours ZS 65.0 69.1 76.6 57.5 76.1 70.8 69.2
Ours w/ PT 66.7 78.3 86.5 78.6 75.9 82.2 78.0
Table 6: Experimental results of Top-5 mAP on the TVSum. † denotes using audio modality. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training, Bold letters indicate the best results.
@@
Ours 89.9 90.5 88.0 87.4 90.0 84.5 77.9 93.2 89.9 80.3 87.1
Ours ZS 44.3 70.9 85.5 65.7 61.1 49.0 44.3 67.3 83.2 48.8 62.0
Ours w/ PT 86.6 74.2 85.9 82.9 90.6 74.5 69.0 92.6 87.6 67.7 81.2

### v2

@@ 4.3.1 Motivation
The QVHighlights dataset contains only 10k training samples. Previous studies [17, 19] have shown that this dataset size is insufficient for training a model to effectively address a complex task like Moment Retrieval. As a result, even training on a noisy pretrain dataset can significantly improve the model’s metrics. For instance, pretraining framework developed in [19] boosted the model performance from 35.47 to 43.63 mAP@avg. Nevertheless, the metrics in the zero-shot mode remained relatively low at only 10.87 mAP@avg [19], leading us to believe that creating a dataset for pretraining remains an unsolved challenge.
@@ 4.3.2 InterVid-MR Dataset
Recent advancements in video-language modeling have been driven by the annotation of large scale, high-quality datasets like InternVid [43]. Unlike VideoCC [27] and other datasets previously used for pretraining [12, 17], InternVid was annotated using LLMs and does not rely on pseudo-labeling, resulting in significantly less annotation noise. To construct our dataset, we used the subset InternVid-10M-FLT [43] derived from the original InternVid.
Unfortunately, InternVid-10M-FLT can not be directly used to train MR models because each caption is linked to only one video segment in the annotations, whereas in practice, multiple video segments may closely match a given query. We have made several steps to address this issue and developed a new dataset named InterVid-MR, which is suitable for MR and HD tasks.
First, we additionally filtered the captions of the InternVid-10M-FLT dataset to remove meaningless ones using ChatGPT  [^2] https://chat.openai.com/chat. Next, to align the data with the QVHighlights dataset, we selected video segments no longer than 150 seconds for each caption from the InternVid-10M-FLT dataset. These segments were created by trimming the original videos around the time intervals associated with the captions.
@@
$S=\text{L2\_Norm}(F_{v})\cdot\text{L2\_Norm}(F_{t})^{T},$ (10)
where $F_{v}$ represents the embedding of the video clip, and $F_{t}$ represents extracted cls embedding of the caption.
@@
After annotating the MR task, we proceeded to annotate the HD task. We used the similarity scores calculated in the previous step to assign each clip within the identified positive intervals a grade ranging from 1 to 4 following the QVHighlights annotation scheme [17]. The grades are calculated using the following formula:
@@
where $S$ is the score from the positive interval to be graded, $\mu$ is the mean score of the positive clips, $\sigma$ is the standard deviation of the positive clips’ scores, and $S_{\text{max}}$ is the maximum score of the positive clips. The constants in Eq. 11 and Eq. 12 were determined empirically. Visualization of the annotation creation process for pretraining is presented in Appendix C.
As a result, we obtain a dataset that can be used to solve both MR and HD tasks simultaneously.
@@ 4.3.3 Impact of the Pretraining
Using the proposed data annotation framework, we collected a dataset consisting of 150k samples. To ensure the effectiveness of the proposed data annotation strategy for solving MR and HD tasks, we measured the zero-shot performance of a model trained on different amounts of pre-train data on the QVHighlights dataset. As shown in Fig. 2, even 20k samples in the training dataset are sufficient to achieve an MR mAP@avg above 45. Using all the data available, we surpassed the 50 MR mAP@avg threshold, which significantly exceeds the results of the pretraining strategy proposed in [19].
3030606090901201201501504545505055556060Number of Pretrain Samples in thousandsmAP@avg (%)QVHighlightsPretrainFinetune
@@ 4.4 Training Details
We train SG-DETR using two attention-pooling layers in the Local Saliency Head and two cross-attention transformer layers in the SGCA Encoder. The model architecture includes three layers of encoders and three layers of decoders. The hidden dimension of the transformer layers is set to 256. We employ the AdamW optimizer [23] with a weight decay of 1e-1. A dropout rate of 0.1 is applied throughout the model, with a higher dropout rate of 0.5 in the projection layers. Our query selector mechanism generates 25 queries per instance. Unless otherwise specified, the following loss balancing parameters are used: $\lambda_{\text{L1}}=10$, $\lambda_{\text{gIoU}}=1$, $\lambda_{\text{CE}}=5$, $\lambda_{\text{Centerness}}=1$, $\lambda_{\text{detr}}=1$, $\lambda_{\text{atss}}=1$, $\lambda_{\text{hl}}=1$, $\lambda_{\text{aux}}=1$. A detailed description of each loss component is provided in Appendix D.
@@
  • QVHighlight: The model is trained on 1xV100 GPU for 150 epochs with a batch size of 128 and a learning rate (lr) of 5e-4. We employ the WarmupMultiStepLR scheduler, utilizing 45 warmup steps, with two milestones set at 100 and 125 epochs. The lr is decayed by a factor of 0.5 at each milestone.
  • Charades-STA: The model is trained on 1xV100 GPU for 100 epochs with a batch size of 64 and a lr of 5e-4. Video features are extracted from 1-second clips, as it was done in [16]. During training, binary highlight scores were generated based on whether a clip overlapped more than halfway with a relevant moment retrieval interval.
  • TACoS: The model is trained on 1xV100 GPU for 120 epochs with a batch size of 32 and a lr of 5e-4. Video features are extracted from 2-second, as it was done in [16]. Binary highlight scores were generated the same way as in Charades-STA.
@@
SG-DETR 72.2 56.6 73.2 55.8 54.1 43.8 69.1
SG-DETR ZS 63.9 49.6 67.5 49.0 48.3 43.0 68.0
SG-DETR w/ PT 74.2 60.4 76.2 60.8 58.8 44.7 71.0
Table 3: Experimental results on the QVHighlights test set. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training. † denotes using audio modality. SG-DETR uses InternVideo2-1b, MR. BLIP uses BLIP-2, and other models use CLIP+SlowFast for FE.
@@
SG-DETR 70.2 49.5 59.1 44.7 29.9 40.9
SG-DETR w/ PT 71.1 52.8 60.7 46.4 33.9 42.4
Table 4: Experimental results on the Charades-STA and TACoS test sets. w/ PT means fine-tuning after pre-training. † denotes using audio modality. SG-DETR uses InternVideo2-1b, MR. BLIP uses BLIP-2, and other models use CLIP+SlowFast for FE.
@@ 4.5 Comparison with the state-of-the-art
We compare SG-DETR against existing state-of-the-art methods on both joint and single-task benchmarks. This quantitative analysis is complemented by a qualitative visualization of our model’s predictions in Appendix B.
@@
To evaluate the effectiveness of SG-DETR in jointly addressing MR and HD tasks, we conducted experiments on the QVHighlights test split, as shown in Tab. 3. Remarkably, even without pretraining, SG-DETR significantly outperforms all existing methods that do not leverage LLMs, and it delivers performance comparable to the MrBLIP model despite the latter having considerably more parameters. After pretraining on our constructed InterVid-MR dataset, SG-DETR achieves competitive zero-shot performance, with 48.3 mAP and 68.0 HIT@1 results. When fine-tuned on the QVHighlights dataset, the pre-trained SG-DETR surpasses all existing methods, including those based on LLMs, with 58.8 mAP and 71.0 HIT@1 results.
We also evaluate on two small-scale highlight detection benchmarks, TVSum [36] and YouTube Highlights [39]. As detailed in Tab. 6 and Tab. 5, our model achieves state-of-the-art results on YouTube Highlights. In contrast, its performance on the data-scarce TVSum is constrained; we attribute this primarily to the dataset’s limited size being insufficient for stable model training. Consequently, fine-tuning with supplemental data does not yield further performance gains on this benchmark.
@@
SG-DETR 75.1 79.7 76.9 71.0 74.5 83.2 76.7
SG-DETR ZS 65.0 69.1 76.6 57.5 76.1 70.8 69.2
SG-DETR w/ PT 66.7 78.3 86.5 78.6 75.9 82.2 78.0
Table 5: Experimental results on the Youtube-HL. The mAP is used as a metric. † denotes using audio modality. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training, bold letters indicate the best results.
@@
SG-DETR 89.9 90.5 88.0 87.4 90.0 84.5 77.9 93.2 89.9 80.3 87.1
SG-DETR ZS 44.3 70.9 85.5 65.7 61.1 49.0 44.3 67.3 83.2 48.8 62.0
SG-DETR w/ PT 86.6 74.2 85.9 82.9 90.6 74.5 69.0 92.6 87.6 67.7 81.2
Table 6: Experimental results on the TVSum. The Top-5 mAP is used as a metric. † denotes using audio modality. w/ PT means fine-tuning after pre-training; ZS means zero-shot inference after pre-training, bold letters indicate the best results.
Moment Retrieval
We use the Charades-STA and TACoS datasets as benchmarks to evaluate the performance of SG-DETR in the moment retrieval task. The results are presented in Table 4. SG-DETR achieves state-of-the-art performance on these benchmarks, both with and without using additional data from our pretraining framework. The results highlight the model’s robustness and its ability to generalize effectively across varying task conditions.
4.6 Ablation study
To evaluate the contribution of each component in the proposed approach, we conducted an ablation study, as shown in Tab. 2. The baseline model (a) utilizes a standard cross-attention transformer for text-to-vision interaction [26], a DETR-based regression head [17, 19], and the local saliency head described in Sec. 3.2. We selected this architecture as our baseline due to its simplicity and reliance on well-established components. In experiment (b), we assess the effect of incorporating the SGCA encoder, detailed in Sec. 3.3. It increases model convergence and pushes performance forward, especially the moment retrieval part. Experiment (c) introduces the IoU scoring mechanism, described in Sec. 3.6, resulting in growth of moment retrieval performance. Next, in experiment (d), we investigate the influence of the GSM module, discussed in Sec. 3.4, on performance. It improves not only MR performance but especially the quality of HD. Finally, in experiment (e), we replace the standard DETR head with a hybrid ATSS-DETR head, outlined in Sec. 3.5, improving MR performance.
5 Conclusions
We introduced SG-DETR, a unified framework for Moment Retrieval and Highlight Detection, and presented an approach for generating pretraining data, enabling us to create the InterVid-MR dataset. Evaluations on benchmarks such as QVHighlights, Charades-STA, and TACoS show that SG-DETR sets state-of-the-art results across all tasks. Ablation studies highlight the effectiveness of key components like the Saliency-Guided Cross-Attention encoder and hybrid ATSS-DETR head, offering a scalable and robust solution for video-language modeling.
6 Limitations
We construct the InterVid-MR dataset using annotations generated by LLMs. Misalignments between captions produced by LLMs and their corresponding video segments may introduce errors in video-text alignment, negatively impacting both model training and performance on fine-grained queries. Additionally, our hyperparameters were selected empirically, and further investigation could potentially yield improved results. Our method also excludes the audio modality, which is a decision motivated not only by modeling considerations but also by the characteristics of the datasets, as QVHighlights and InterVid-MR primarily focus on visual content and lack aligned audio annotations.
Supplementary Material
MR HD
Attention Type R1 mAP $\geq$ Very Good
@0.5 @0.7 @0.5 @0.75 Avg. mAP HIT@1
(a) CA 70.1 $\pm$ 0.3 53.7 $\pm$ 0.5 70.0 $\pm$ 0.1 49.3 $\pm$ 0.2 48.3 $\pm$ 0.4 41.9 $\pm$ 0.1 68.6 $\pm$ 0.8
(b) ACA 67.7 $\pm$ 0.6 51.5 $\pm$ 0.7 68.22 $\pm$ 0.2 47.7 $\pm$ 0.8 47.1 $\pm$ 0.7 42.0 $\pm$ 0.1 69.41 $\pm$ 0.5
(c) SGCA 71.1 $\pm$ 0.5 56.5 $\pm$ 0.8 71.0 $\pm$ 0.4 52.2 $\pm$ 0.7 50.1 $\pm$ 0.6 42.2 $\pm$ 0.1 70.1 $\pm$ 0.6
Table 7: Ablation study of attention types on QVHighlights val split. CA, ACA, and SGCA stand for Cross Attention, Adaptive Cross-Attention, and Saliency-Guided Cross-Attention. Results are reported as mean $\pm$ standard deviation, averaged over three runs.
Appendix A SG-Attention: Computational Efficiency
SG-Attention only minimally increases computational cost compared to standard dot-product attention. In the conventional setup, the main cost arises from two matrix multiplications: $QK^{\top}$ and $\mathrm{AttnWeights}\times V$. Each multiplication requires $2L^{2}E$ multiply-add operations, leading to a total of $4L^{2}E$. By contrast, SG-Attention performs one additional elementwise multiplication of size $L^{2}$, which is negligible when $E$ is large. For instance, with $L=76$ and $E=256$, standard attention incurs approximately $5.915\times 10^{6}$ operations, whereas SG-Attention adds only about $5.776\times 10^{3}$ operations (less than $1\%$ overhead). Consequently, this saliency mechanism imposes only a minor computational burden and does not impede real-time performance.
Appendix B Comparative Study of Attention Mechanisms
We evaluate and compare three distinct attention mechanisms for fusing textual information into video feature representations: standard Cross-Attention, the Adaptive Cross-Attention proposed by CGDETR [25], and our novel Saliency-Guided Attention. The results for the baseline architecture, trained on the QVHighlights dataset with each of the three attention variants, are summarized in Tab. 7.
Appendix C Training Objectives
Our model is trained using losses categorized into three main groups.
Highlight Detection Task
Moment Retrieval Task
For the MR task, we use CE loss, generalized IoU (GIoU) loss [32], and smooth L1 loss to train both DETR and ATSS detection heads. For the ATSS head [48], we also incorporate Centerness Loss [40] to enhance localization precision. In the DETR head, CE loss is applied to IoU head predictions. Losses for auxiliary DETR queries are computed similarly to primary queries. The objectives for the task are:
Auxiliary Losses
$\mathcal{L}_{aux}=\mathcal{L}_{bce}+\mathcal{L}_{align}$ (16)
Overall Objective
Appendix D Qualitative Results and Analysis
In Fig. 3, we present visualizations of predictions made by the models on the QVHighlights validation dataset. Compared to existing methods such as CG-DETR [25] and TR-DETR [37], SG-DETR achieves more precise and coherent highlight detection results, as evidenced by improvements in both retrieval accuracy and highlight score distributions.
Appendix E Visualization of the Pretrain Annotation Framework
To clearly illustrate how the InterVid-MR pretraining dataset was created, we present key statistics related to the annotation generation process in Fig. 4. Although the annotation process is straightforward, the visualization confirms that the resulting annotations achieve high quality.
Figure 3: Qualitative results
Figure 4: Design of the Preatrain Annotation